OpenStudy (anonymous):

simplify the rational expression. State any restrictions on the variable. q^2 + 11q +24/q^2 - 5q - 24

OpenStudy (anonymous):

@tcarroll010

OpenStudy (abb0t):

It might help to simplify by factoring the numerator and denominator and cross cancelling.

OpenStudy (abb0t):

\[\frac{ (x+3)(x+8) }{ (x+3)(x-8) } = \]

OpenStudy (anonymous):

\[\frac{ (q + 8)\cancel {(q + 3)} }{ (q - 8)\cancel {(q + 3)} } = \frac{ q + 8 }{ q - 8 }\]and you cannot have q =8 or q = -3

OpenStudy (abb0t):

How did you do that crossing out?! :O

OpenStudy (anonymous):

First, good luck in all of youstudies and thx for the recognition! Second, this box uses Latex formatting, so I did a backslash cancel and put what I wanted cancelled in between curly braces. Try it!

OpenStudy (anonymous):

I spelled out backslash but when you use it, you just use the symbol, then the word cancel, then followed by what you want to cancel inside curly braces.